OpenStudy (anonymous):

Solve the following system using the elimination method: 6x - 2y = 4 -2x + 4y = 12

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

hi c:

OpenStudy (anonymous):

hey!

zepdrix (zepdrix):

To eliminate, we need to have the `same` coefficient on one of our variables.

zepdrix (zepdrix):

To to match up our x's, we would want a 6 in front of each. How could we fix up the second equation to make that happen?

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

not sure

zepdrix (zepdrix):

Hmm it currently has a value of -2. Let's multiply it the second equation by 3 and see what happens. (We're multiplying both sides by 3, otherwise things become unbalanced).\[\Large\rm 3(-2x + 4y) = (12)3\]Distributing the 3 gives us,\[\Large\rm -6x + 12y =36\]

zepdrix (zepdrix):

Now our system is,\[\Large\rm 6x - 2y = 4\]\[\Large\rm -6x + 12y =36\]

zepdrix (zepdrix):

The reason we're doing this is so that we can either add or subtract our equations together, and hopefully one of the variables will cancel out.

OpenStudy (anonymous):

(0,3) (1, 7/2)

OpenStudy (anonymous):

inconsistent

zepdrix (zepdrix):

? =o

OpenStudy (anonymous):

I calculated it Im wrong I guess??

zepdrix (zepdrix):

I don't know what you did...

zepdrix (zepdrix):

Since our equations have `opposite signs` on the x's, we'll want to `add` them together.

zepdrix (zepdrix):

So 6x and -6x add to zero. And we're left with:\[\Large\rm 10y=40\]

OpenStudy (anonymous):

4

zepdrix (zepdrix):